OpenStudy (anonymous):

Evaluate the following using l'Hospital's Rule: 1) lim x->∞ ln(√x)/x^2 2)lim x-> 1 (3-2x)^(1/x-1)

OpenStudy (turingtest):

you know l'Hospital's rule? just take the derivative of the top and bottom when you get an indeterminate form like this and then take the limit again if that is still indeterminate, use l'Hospital again until it works looks to me like you will have to exponentiate the 2nd one

OpenStudy (anonymous):

I tried, I keep on ending up up with 0/0 for the first one.

OpenStudy (turingtest):

what is the derivative of\[\ln\sqrt x\]?

OpenStudy (anonymous):

1/(√x)

OpenStudy (turingtest):

don't forget to use the chain rule...

OpenStudy (turingtest):

or you could simplify the log first, makes it easier to see

OpenStudy (turingtest):

\[\ln\sqrt x=\frac12\ln x\]

OpenStudy (anonymous):

Okay, this is what I got using the chain rule, 1/0.5X^(-0.5).0.25X^(-0.5)

OpenStudy (anonymous):

Is that okay?

OpenStudy (turingtest):

erm, I'm trying to tell, but I don't see where you got 0.25

OpenStudy (anonymous):

Okay, I think Ill just try it again with the 1/2lnx then.

OpenStudy (turingtest):

why not just note that you can simplify the logarithm as I did above that should at least make it clear what answer you should get if you do the chain rule right

OpenStudy (turingtest):

yeah, safer bet, but just to show you that the chain rule works...

OpenStudy (turingtest):

the chain rule on\[\frac d{dx}\ln\sqrt x\]would give\[x^{-1/2}\cdot\frac12x^{-1/2}=\text{(same thing, you write it)}\]

OpenStudy (anonymous):

Okay, thanks.

OpenStudy (turingtest):

so what is the derivative of the numerator?

OpenStudy (anonymous):

its 1/2. 1/x = -2/x...I think.

OpenStudy (turingtest):

\[\frac d{dx}\ln\sqrt x=\frac12\frac d{dx}\ln x=\frac12\cdot\frac1x=\frac1{2x}\]what is the derivative of the denominator?

OpenStudy (anonymous):

But -2x works too, right?

OpenStudy (turingtest):

no, why negative?

OpenStudy (turingtest):

and x can't be in the numerator

OpenStudy (turingtest):

...of the numerator that is :P

OpenStudy (anonymous):

Oh okay, my bad.

OpenStudy (turingtest):

\[\ln\sqrt x=\ln x^{1/2}=\frac12\ln x\]and the derivative of lnx is 1/x, so the derivative of the above is 1/(2x)

OpenStudy (turingtest):

that's the numerator, what about the denominator?

OpenStudy (anonymous):

2x, I know that one :)

OpenStudy (turingtest):

good, so what is the limit now?

OpenStudy (anonymous):

so its 1/2x/2x

OpenStudy (turingtest):

which simplifies to what?

OpenStudy (anonymous):

1/4X^2 im sorry im really bad at this

OpenStudy (turingtest):

no, you got it right :) so now we can take the limit, and it is....?

OpenStudy (anonymous):

1/4(0)^2 = 1/0 = DNE?

OpenStudy (turingtest):

what is x approaching? I don't think it's zero, check your post

OpenStudy (anonymous):

or = infinity is it?

OpenStudy (anonymous):

Its approaching infinity, so don't you substitute x with 0?

OpenStudy (turingtest):

no, you would substitute x with zero when x approaches zero you have confused the issue a bit...

OpenStudy (turingtest):

\[\lim_{x\to\infty}\frac1x=?\]as x goes to infinity, the \(fraction\) goes to zero why? because as we put in larger and larger values of x the fraction gets smaller and smaller indefinitely.

OpenStudy (anonymous):

Okay..

OpenStudy (turingtest):

so as x gets very large in your expression\[\lim_{x\to\infty}\frac1{4x^2}\]what number does the fraction \(\large\frac1{4x^2}\) approach ?

OpenStudy (anonymous):

1? I'm so confused.

OpenStudy (turingtest):

x is getting larger and larger what is 1/4x^2 when x=1? when x=2 ? when x=10 ? when x=100 ? try plugging in the numbers if you don't see what is happening i.e. as x gets bigger what happens to 1/4x^2 ? what number is it getting close to?

OpenStudy (anonymous):

Is it approaching zero?

OpenStudy (turingtest):

yes :)

OpenStudy (turingtest):

to write an ill-formed expression, it can be said that\[\lim_{x\to\infty}\frac1x=\frac1\infty=0\]because anything divided by infinity is zero. better to see the concept though, what Is wrote is less than formal notation I think...

OpenStudy (turingtest):

the same deal is happening with your expression, or any where we wind up with \(x^n\) where \(n\ge1\) in the denominator: as x gets big, 1/x^n gets infinitely small and approaches zero

OpenStudy (anonymous):

Okay, makes sense but what would I substitute x with in my working?

OpenStudy (turingtest):

well, you could use the informal notation I suggest, but by the time I do problems with l'Hospital it seems sufficient to just put \[\lim_{x\to\infty}\frac1x=0\]and not have to give a big explanation as to why the limit is fairly obvious once you grasp the concept however @Zarkon may have a better suggestion as to what a teacher would want to see you write

OpenStudy (anonymous):

Okay.

OpenStudy (turingtest):

here is the best I could muster if you wanted me to write something out\[\lim_{x\to\infty}\frac1{4x^2}=\frac14\lim_{x\to\infty}\frac1{x^2}=\frac14\cdot\frac1\infty=\frac1\infty=0\]which is a big ridiculous song and dance really. Infinity isn't even a number so writing = is sorta... well here comes Zarkon to explain

OpenStudy (zarkon):

\[\lim_{x\to\infty}\frac1x=0\] is usually given as a small theorem in most intro calculus texts.

OpenStudy (turingtest):

good let it rest there then :)

OpenStudy (anonymous):

Thank you, guys! :)

OpenStudy (zarkon):

some even give \[\lim_{x\to\infty}\frac{1}{x^r}=0\] for \[r>0\]

OpenStudy (turingtest):

oh I wrote \(r\ge1\) above, my bad :(

OpenStudy (anonymous):

It's fine. Think you could help me with the second one? :) pls

OpenStudy (turingtest):

I think you should post it separately, I have a feeling you will need a fair amount of explanation for that one. Also I have to leave soon, so hopefully someone else can take the reigns on this. basically you will use exponents to transform the limit, which if the process is unfamiliar to you will take some time to clarify

OpenStudy (turingtest):

after doing the exponent trick you can apply l'hospital pretty easily

OpenStudy (anonymous):

Okay, I'll do that. Thank you so much! Really appreciate it.

OpenStudy (turingtest):

the exponent trick I refer to is this\[\lim_{x\to a}f(x)=\exp\left(\lim_{x\to a}[\ln f(x)]\right)\]you're welcome, good luck!